[S]             v (nM)      (nM s-1) 1.25          1.10 3.60          4.15 5.83&

[S]             v

(nM)      (nM s-1)

1.25          1.10

3.60          4.15

5.83         6.54

7.34        9.38

9.12        11.71

Calculate the Vmax and Km values using the following methods with work.

Line weaver Burke:

Hanes Woolf:

Eadie Hofstee:
Answer #1

Lineweaver Burk:

Equation:

1 V Km 1 SV mar Km

S V 1/S 1/V
1.25 1.1 0.8 0.909091
3.6 4.15 0.277778 0.240964
5.83 6.54 0.171527 0.152905
7.34 9.38 0.13624 0.10661
9.12 11.71 0.109649 0.085397

Lineweaver Burk Plot = 1.203x0.061 R=0.997 NT TIITTO 0.2 0.4 0.6 0.8 1/[S]

Comparing with the equation in the plot the standard equation

1/Km = 0.061

=> Km =16.39

Km/Vmax = 1.203

=> Vmax=16.39/1.203

=> Vmax = 13.62 nM/s

Hanes Woolf:

Equation

(s) _ Is Km V Vmar. Vmar

V S S/V
1.1 1.25 1.136364
4.15 3.6 0.86747
6.54 5.83 0.891437
9.38 7.34 0.782516
11.71 9.12 0.778822

Hanes Woolf Plot y=-0.041x1118 R = 0.789 NIS) O 2 4 6 8 10 [s]

Comparing the equation in the plot with the standard equation

1/Vmax=0.041

=>Vmax= 24.39 nM/s

Km/Vmax = 1.118

=> Km = 24.39*1.118

=>Km=27.268

Eadie Hofstee:

Equation:

V = -Km * V - + V mar 1s

S V/S V
1.25 0.88 1.1
3.6 1.152778 4.15
5.83 1.121784 6.54
7.34 1.277929 9.38
9.12 1.283991 11.71

Eadie Hofstee Plot y= 23.37x20.14 R2 = 0.840 0 0.2 0.4 0.6 1 1.2 1.4 0.8 v/[S]

Comparing with the equation in the plot the standard equation

Vmax= 20.14 nM/s

Km= 23.37

Summary of all answer:

Lineweaver Burk Hanes Woolf Eadie Hofstee
Vmax 13.62 24.39 20.14
Km 16.39 27.268 23.37